OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y – 3z = –3 x – 2y – 2z = –3 x + y + 3z = –2

OpenStudy (anonymous):

i hate these they take forever

OpenStudy (anonymous):

method is long and tedious, eliminate one variable and then substitute back it is much easier to cheat http://www.wolframalpha.com/input/?i=4x+%E2%80%93+y+%E2%80%93+3z+%3D+%E2%80%933%2C+x+%E2%80%93+2y+%E2%80%93+2z+%3D+%E2%80%933+%2Cx+%2B+y+%2B+3z+%3D+%E2%80%932

OpenStudy (anonymous):

Yeah i just don't get it all...

OpenStudy (anonymous):

oh no that is wrong

OpenStudy (anonymous):

subtract third equation from second get \[-3y-5z=-1\] or \[3y+5z=1\]

OpenStudy (anonymous):

then multiply the second equation by -4 and add to the first \[4x – y – 3z = –3 \] \[-4x-4 y -12z = 8\] \[-5y-15z=5\] or \[y+3z=-1\]

OpenStudy (anonymous):

now we have two equation \[y+3z=-1\] \[3y+5z=1\] now it is easier to solve you get \(y=2,z=-1\) and substitute back to solve for \(x\)

OpenStudy (anonymous):

You will get x=-1